OpenStudy (anonymous):

Jose received $30 for his birthday and put it in his piggy bank. Each week, he puts 4 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 4w + 30 where w is the number of weeks he has saved 4 dollars. His sister Marisol makes a table of reasonable values for w. Marisol selects 16 as a value for w. What is the corresponding table entry for D? a. 34 b. 434 c. 50 d. 94

OpenStudy (dayakar):

d

OpenStudy (anonymous):

are you sure it is d

OpenStudy (texaschic101):

sub in 16 for w... D = 4(16) + 30 D = 64 + 30 D = 94

OpenStudy (anonymous):

Step 1: Substitute D=4(16)+30 Step 2: Simplify D= 64+30 Step 3: Simplify D=94 Step 4: Answer D=94 choice letter d

OpenStudy (anonymous):

thanks

OpenStudy (dayakar):

d